I was five times as old as my sister five years ago, and I am three times older than her now. How old am I?
Answer by scott8148(6628)

then ___ x - 5 = 5 (s - 5) ___ x = 5s - 20

now ___ x = 3s

substituting ___ 3s = 5s - 20 ___ s = 10

substituting ___ x = 3(10) = 30